OpenStudy (brucebaner):

Relationship B has a greater rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose all answers that are correct.

OpenStudy (brucebaner):

A y = 3/4x B y= 0.6x C y =2/3x D y = 1/4x

OpenStudy (brucebaner):

@KamiBug

OpenStudy (brucebaner):

@keke_luvs_u

OpenStudy (brucebaner):

@sangya21

OpenStudy (anonymous):

Find the slope of graph of Relationship A then find equation of B B's slope is greater than A's

OpenStudy (brucebaner):

@geerky42